# Image Generation from Contextually-Contradictory Prompts

### v1

@@ Appendix A Appendix
Table 4. Full LLM prompt instruction for time step dependent prompt conditioning, used to decompose prompts by denoising stage.
@@
- The length of switch_ prompts_steps should be one less than prompts_list. - Do not include any text outside this structure.
A.1. LLM Instruction for Prompt Decomposition
Tables 4 and 5 detail the full LLM instruction used for our method’s decomposition, along with the corresponding in-context examples. In a single inference pass, our method detects contextual contradictions, generates proxy prompts, and assigns timestep intervals.
Table 5. In-context examples for time step dependent prompt conditioning, illustrating reasoning steps and prompt decomposition across denoising stages.
@@ A.1. LLM Instruction for Prompt Decomposition
A.2. VLM evaluation
We evaluate text-to-image alignment for the different models on the benchmarks using the VLM GPT-4o. The instruction prompt used for the VLM is presented in Table 6.
A.3. ContraBench
To evaluate image generation under prompts containing contextual contradictions, we introduce ContraBench, presented in Table 7. This dataset consists of 40 prompts generated using ChatGPT, each intentionally constructed to include visual or semantic inconsistencies challenge standard diffusion models. The prompts span a range of contradiction types, including physical implausibility, role violation, object misplacement, and surreal transformations. ContraBench serves as a targeted benchmark for assessing a model’s ability to produce visually coherent and semantically plausible images in the presence of contradictory concepts.
A.4. Method Generalization to Different Models
Our method’s proxy prompt injection mechanism integrates seamlessly into existing inference pipelines without requiring any architectural modifications, and it is compatible with a range of pretrained diffusion models. Differences in denoising schedules across models are handled entirely during the prompt decomposition stage. See Figure 11 for an example of how our method integrates with SD3 without any modifications.
@@ A.4. Method Generalization to Different Models
Table 6. VLM instruction for prompt alignment. Instructions provided to the VLM for scoring semantic alignment of generated images.
@@
SD3 Ours(SD3) SD3 Ours(SD3)
Refer to caption Refer to caption Refer to caption Refer to caption
“A grown man has a baby’s pacifier in his mouth.” “A camping tent is inside a bed room.”
Refer to caption Refer to caption Refer to caption Refer to caption
“A mouse eats a snake.” “A dragon is blowing water.”
Figure 11. Generated images from our method applied on SD3.
Table 7. ContraBench prompt set.
ID Prompt
1 A professional boxer does a split
2 A bear performing a handstand in the park
3 A photorealistic photo of SpongeBob SquarePants dancing ballet
4 A cowboy swimming competitively in an Olympic pool
5 A cruise ship parked in a bathtub
6 A man giving a piggyback ride to an elephant
7 A zebra climbing a tree
8 A coffee machine dispensing glitter
9 A vending machine in a human running posture
10 A ballerina aggressively flipping a table
11 A bathtub floating above a desert in a tornado
12 A monkey juggles tiny elephants
13 a woman has a marine haircut
14 A tower with two hands
15 An archer is shooting flowers with a bow
16 A baseball player backswing a yellow ball with a golf club
17 A barn built atop a skyscraper rooftop
18 A cat balancing a skyscraper on its nose
19 A cow grazing on a city rooftop
20 A fireplace burning inside an igloo
21 A mosquito pulling a royal carriage through Times Square
22 A grandma is ice skating on the roof
23 A muscular ferret in the woods
24 A house with a circular door
25 A photorealistic image of a bear ironing clothes in a laundry room
26 A pizza being used as an umbrella in the rain
27 A cubist lion hiding in a photorealistic jungle
28 A chicken is smiling
29 A realistic photo of an elephant wearing slippers
30 A computer mouse eating a piece of cheese
31 A horse taking a selfie with a smartphone
32 A sheep practicing yoga on a mat
33 a snake eating a small golden guitar
34 A soccer field painted on a grain of rice
35 A snake with feet
36 A woman brushing her teeth with a paintbrush
37 A horse with a hump
38 A hyperrealistic unicorn made of origami
39 A library printed on a butterfly’s wings
40 A bodybuilder balancing on pointe shoes
Table 8. Examples of prompt decomposition for contradictory prompts using our method.
Original Prompt Explanation Prompt Sequence Switch Step(s)
A grown man has a baby’s pacifier in his mouth. The grown man is established first for structure, while the pacifier, being a small but semantically significant detail, is added once the main identity is resolved. (1) A grown man with a small object in his mouth (2) A grown man has a baby’s pacifier in his mouth 4
A dragon is blowing water. Dragons are more commonly depicted blowing fire. A proxy of white smoke is visually similar to water mist in texture, stabilizing the emission process before resolving the surreal water emission. (1) A dragon blowing white smoke (2) A dragon blowing water 3
A pizza with grape toppings. Pizza with traditional toppings stabilizes the geometry and color before introducing the visually similar yet unusual grape topping. (1) A pizza with pepperoni toppings (2) A pizza with grape toppings 3
A coin floats on the surface of the water. Coins typically sink in water, not float. Starting with a leaf—an object that naturally floats—ensures that this behavior within the scene is handled correctly before introducing the coin. (1) A leaf floats on the surface of the water (2) A coin floats on the surface of the water 4
A cockatoo parrot swimming in the ocean. Cockatoos are birds and naturally do not swim; starting with a simple bird on water stabilizes position and motion. Progressing to a duck, before introducing the cockatoo parrot, eases the transition into the final surreal visual. (1) A duck swimming in the ocean (2) A parrot swimming in the ocean (3) A cockatoo parrot swimming in the ocean 3, 6
A woman writing with a dart. Writing is typically associated with a pen or pencil. Starting with a pen stabilizes the action before introducing the unusual tool. (1) A woman writing with a pen (2) A woman writing with a dart 3
Shrek is blue. Shrek is a distinct character with a recognizable green color. Using a simple b̈lue ogreïnitially sets the stage for a color change before fully introducing Shrek to ensure visual coherence. (1) A blue ogre (2) Shrek is blue 3

### v2

@@ 11 VLM Evaluation
Table 8: Full LLM prompt instruction SAP, used to decompose prompts by denoising stages.
@@
- The length of switch_ prompts_steps should be one less than prompts_list.
- Do not include any text outside this structure.
Table 9: In-context examples SAP, illustrating reasoning steps and prompt decomposition across denoising stages.
@@
Table 10: Prompt decompositions of contradictory prompts using SAP.
Original Prompt Explanation Prompt Sequence Switch Step(s)
A grown man has a baby’s pacifier in his mouth. The grown man is established first for structure, while the pacifier, being a small but semantically significant detail, is added once the main identity is resolved. 1. A grown man with a small object in his mouth 2. A grown man has a baby’s pacifier in his mouth 4
A dragon is blowing water. Dragons are more commonly depicted blowing fire. A proxy of white smoke is visually similar to water mist in texture, stabilizing the emission process before resolving the surreal water emission. 1. A dragon blowing white smoke 2. A dragon blowing water 3
A pizza with grape toppings. Pizza with traditional toppings stabilizes the geometry and color before introducing the visually similar yet unusual grape topping. 1. A pizza with pepperoni toppings 2. A pizza with grape toppings 3
A coin floats on the surface of the water. Coins typically sink in water, not float. Starting with a leaf—an object that naturally floats—ensures that this behavior within the scene is handled correctly before introducing the coin. 1. A leaf floats on the surface of the water 2. A coin floats on the surface of the water 4
A cockatoo parrot swimming in the ocean. Cockatoos are birds and naturally do not swim; starting with a simple bird on water stabilizes position and motion. Progressing to a duck, before introducing the cockatoo parrot, eases the transition into the final surreal visual. 1. A duck swimming in the ocean 2. A parrot swimming in the ocean 3. A cockatoo parrot swimming in the ocean 3, 6
Shrek is blue. Shrek is a distinct character with a recognizable green color. Using a simple b̈lue ogreïnitially sets the stage for a color change before fully introducing Shrek to ensure visual coherence. 1. A blue ogre 2. Shrek is blue 3
A professional boxer does a split. Professional boxers are typically shown in athletic stances related to fighting, not performing a split. Starting with a gymnast performing a split supports the action, introducing a boxer in similar attire balances identity shift without disrupting the pose. 1. A gymnast performing a split 2. A boxer performing a split 3. A professional boxer doing a split 3, 6
Table 11: ContraBench. A curated benchmark of 40 contradictory prompts for evaluating text-to-image models.
ID Prompt ID Prompt
1 A professional boxer does a split 21 A mosquito pulling a royal carriage through Times Square
2 A bear performing a handstand in the park 22 A grandma is ice skating on the roof
3 A bodybuilder balancing on point shoes 23 A baseball player backswing a yellow ball with a golf club
4 A chicken is smiling 24 A house with a circular door
5 A cruise ship parked in a bathtub 25 A photorealistic image of a bear ironing clothes in a laundry room
6 A man giving a piggyback ride to an elephant 26 A pizza being used as an umbrella in the rain
7 A zebra climbing a tree 27 A cubist lion hiding in a photorealistic jungle
8 A coffee machine dispensing glitter 28 A cowboy swimming competitively in an Olympic pool
9 A vending machine in a human running posture 29 A realistic photo of an elephant wearing slippers
10 A ballerina aggressively flipping a table 30 A computer mouse eating a piece of cheese
11 A bathtub floating above a desert in a tornado 31 A horse taking a selfie with a smartphone
12 A monkey juggles tiny elephants 32 A sheep practicing yoga on a mat
13 A woman has a marine haircut 33 A snake eating a small golden guitar
14 A tower with two hands 34 A soccer field painted on a grain of rice
15 An archer is shooting flowers with a bow 35 A snake with feet
16 A muscular ferret in the woods 36 A woman brushing her teeth with a paintbrush
17 A barn built atop a skyscraper rooftop 37 A horse with a hump
18 A cat balancing a skyscraper on its nose 38 A hyperrealistic unicorn made of origami
19 A cow grazing on a city rooftop 39 A library printed on a butterfly’s wings
20 A fireplace burning inside an igloo 40 A photorealistic photo of SpongeBob SquarePants dancing ballet
Table 12: Whoops-Hard. A curated subset of 100 challenging prompts from the Whoops! benchmark.
ID Prompt ID Prompt
1 A bouquet of flowers is upside down in a vase 51 A Japanese tea ceremony uses coffee instead of tea
2 A man is welding without a mask 52 A wagon is being pushed from behind by two opposite facing horses
3 A man eats hamburgers in a baby chair 53 The Girl with a Pearl Earring wears a golden hoop earring
4 A turn right street sign with a left turn arrow 54 A chandelier is hanging low to the ground
5 Goldilocks sleeps with four bears 55 The portrait of the Mona Lisa depicts a stern male face
6 A cake wishes a happy 202nd birthday 56 A car with the steering wheel right in the middle of the dashboard
7 Children are unhappy at Disneyland 57 A pagoda sits in front of the Eiffel Tower
8 An orange carved as a Jack O’Lantern 58 A man without protection next to a swarm of bees
9 A pen is being sharpened in a pencil sharpener 59 A kiwi bird in a green bamboo forest
10 Steve Jobs demonstrating a Microsoft tablet 60 The Sphinx is decorated like a sarcophagus outside a Mayan temple
11 Shrek is blue 61 A butterfly is in a bee’s hive
12 A MacBook with a pear logo on it 62 A rainbow colored tank
13 A woman hits an eight ball with a racket 63 Movie goers nibble on vegetables instead of popcorn
14 Vikings ride on public transportation 64 A grown man has a baby’s pacifier in his mouth
15 A gift wrapped junked car 65 A full pepper shaker turned upside down with nothing coming out
16 A rainbow is filling the stormy sky at night 66 The Tiger King, Joe Exotic, poses with an adult saber-tooth tiger
17 John Lennon using a MacBook 67 A scale is balanced with one side full and the other empty
18 Michelangelo’s David is covered by a fig leaf 68 A pizza box is full of sushi
19 Chuck Norris struggles to lift weights 69 A man wearing a dog recovery cone collar while staring at his dog
20 Paratroopers deploy out of hot air balloons 70 A woman’s mirror reflection is wearing different clothes
21 A train on asphalt 71 A woman using an umbrella made of fishnet in the rain
22 Lionel Messi playing tennis 72 A field of sunflowers with pink petals
23 A man jumping into an empty swimming pool 73 An eagle swimming under water
24 An airplane inside a small car garage 74 A woman stands in front of a reversed reflection in a mirror
25 An upside down knife about to slice a tomato 75 Stars visible in the sky with a bright afternoon sun
26 Dirty dishes in a bathroom sink 76 A car with an upside down Mercedes-Benz logo
27 A roulette wheel used as a dart board 77 An owl perched upside down on a branch
28 A smartphone plugged into a typewriter 78 A man in a wheelchair ascends steps
29 A passenger plane parked in a parking lot 79 Bach using sound mixing equipment
30 Guests are laughing at a funeral 80 A steam train on a track twisted like a roller coaster
31 A cat chasing a dog down the street 81 Roman centurions fire a cannon
32 The Statue of Liberty is holding a sword 82 A crab with four claws
33 A Rubik’s cube with ten purple squares 83 Elon Musk wearing a shirt with a Meta logo
34 A girl roller skating on an ice rink 84 A compass with North South South West points
35 A butterfly swimming under the ocean 85 A glass carafe upside down with contents not pouring
36 Lightning striking a shack on a sunny day 86 Princess Diana standing in front of her grown son, Prince Harry
37 The Cookie Monster is eating apples 87 A children’s playground set in the color black
38 A man is given a purple blood transfusion 88 A mug of hot tea with a plastic straw
39 An unpeeled banana in a blender 89 A whole pomegranate inside a corked glass bottle
40 A square apple 90 Belle from Beauty and the Beast about to kiss the Frog Prince
41 A place setting has two knives 91 A person’s feet facing opposite directions
42 A koala in an Asian landscape 92 A bowl of cereal in water
43 A mouse eats a snake 93 A boy playing frisbee with a porcelain disk
44 A field of carrots growing above ground 94 A chef prepares a painting
45 A pregnant woman eating raw salmon 95 A dragon blowing water
46 A tiger staring at zebras in the savanna 96 The lip of a pitcher on the same side as the handle
47 Albert Einstein driving a drag racing car 97 Greta Thunberg holding a disposable plastic cup
48 A soccer player about to kick a bowling ball 98 A fortune teller predicting the future with a basketball
49 An old man riding a unicycle 99 A balloon lifting up a package
50 A hockey player drives a golf ball down the ice 100 Bruce Lee in a yellow leotard and tutu practicing ballet
@@ 49 An old man riding a unicycle 99 A balloon lifting up a package
Table 13: VLM instruction for evaluation. Used by GPT-4o to score semantic alignment of generated images.